Derive the moment of inertia of a rectangular plate having sides a and b about an axis passing through one end and perpendicular to the plate.

Grade:12

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

I = integral (r^2) *dm
rho = m/ab
dm = rho*b*dr
therefore, dI = r^2 * rho*b*dr. plugging in rho we get
dI = (r^2 * m *b*dr )/ab

If you integrate this from 0 to a, you get
I = (m*a^2)/3

Using perpendicular axis theorem, Iz = Ix + Iy so
Iz = [m (a^2 + b^2)]/3

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi